# The DeepSpeak Dataset

### v3

@@ Abstract
We describe a large-scale dataset – DeepSpeak – of real and deepfake footage of people talking and gesturing in front of their webcams. The real videos in this dataset consist of a total of 50 hours of footage from 500 diverse individuals. Constituting more than 50 hours of footage, the fake videos consist of a range of different state-of-the-art avatar, face-swap, and lip-sync deepfakes with natural and AI-generated voices. We are regularly releasing updated versions of this dataset with the latest deepfake technologies. This preprint describes the construction of versions 1.0, 1.1, and 2.0. This dataset is made freely available for research and non-commercial uses; requests for commercial use will be considered.
Keywords Deepfakes  $\cdot$ Generative AI  $\cdot$ Dataset  $\cdot$ Digital Forensics
@@ 1 Introduction
Large-scale and diverse datasets are critical to the academic literature to aid in both system development and evaluation. We describe an audio and video dataset designed to aid the digital-forensic community [^1] We named our dataset DeepSpeak before the release of the DeepSeek AI model. This dataset consists of 100 hours of real and deepfake video of people talking and gesturing. The real videos were recorded with consent from the participants. The deepfakes consist of avatar (from three generators), face-swap deepfakes (with multiple variants from three generators), lip-sync deepfakes (from four generators), and audio deepfakes (from three generators) spliced into a subset of the lip-sync deepfakes. Shown in Table 1 is a comparison of the characteristics of the different versions of our dataset as compared to recent (past seven years) datasets.
We describe the details of the collection of the real video and how the various deepfake audio and video are created. Each version of the dataset is available by request at the following locations:
Given the pace at which the underlying deepfake technology is progressing, it is critical that evaluation datasets keep up with the latest technologies. To this end, we expect to release an updated dataset once to twice a year, expanding all aspects of the dataset.
@@
FaceForensics  [25] 2018 NA 1,004 N $\checkmark$ - - - 2,008
FaceForensics++ [26] 2019 NA 1,000 N $\checkmark$ - $\checkmark$ - 4,000
DFDC [6] 2020 3,426 23,654 Y $\checkmark$ - - - 104,500
Celeb-DF [15] 2020 59 590 N $\checkmark$ - - - 5,639
WildDeepfake [43] 2020 707 3,805 N NA NA NA NA 3,509
DeeperForensics [11] 2020 100 50,000 Y $\checkmark$ - - - 10,000
FakeAVCeleb [12] 2021 600 570 N $\checkmark$ $\checkmark$ - - 25,000
ForgeryNet [10] 2021 5,400 99,630 Y/N $\checkmark$ - - - 121,617
LAV-DF [2] 2022 153 36,431 N - - $\checkmark$ - 99,873
DF40 [35] 2024 NA NA N - 100,000+
DeepSpeak v1.0 2024 219 6,226 Y $\checkmark$ $\checkmark$ - $\checkmark$ 6,799
DeepSpeak v1.1 2025 220 6,667 Y $\checkmark$ $\checkmark$ - $\checkmark$ 6,796
DeepSpeak v2.0 2025 280 9,376 Y $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 7,209
@@ 2 Data Collection
Participants were crowd-sourced through the Prolific research recruitment platform. Participants were asked to give their consent for including their recordings, without any other identifying information, in a public dataset [^2] The precise consent language was: “This dataset will be used for research purposes for detecting deepfakes. Please note that your recordings will be made public in a dataset, but no other identifying information will be shared outside of our research group.”. Participants were selected from a stratified sample ensuring equal distribution of gender, with all participants being native English speakers and U.S. residents.
Each participant completed a survey in which they were instructed to record themselves responding to 32 (version 1.0) and 35 (version 2.0) separate prompts (see Appendix A). Participants were paid \$7 for their time. The first two prompts were used for voice-clone training data (see Section 3). The remaining prompts were divided into four categories: (1) standardized scripted responses in which each participant read the same prompt; (2) randomized scripted responses in which participants read a randomized prompt; (3) unscripted responses in which participants responded to questions; and (4) actions in which participants performed simple actions.
Both audio and video were recorded using a custom-built Google Chrome web application. Recordings were captured as .webm files with a bitrate of 8 megabits per second, using the Google VP9 codec for video compression. The target resolution was set at 1280 $\times$× 720 pixels, but users with limited bandwidth were able to record at lower resolution of 640 $\times$× 480 pixels. The JavaScript and Python repository for this web application is available at https://github.com/sbarrington/UCB-web-video-recorder-app-main.
Participants were given written and visual instructions to allow them to practice recording themselves and test their hardware. Participants were asked to adhere to a series of recording conditions intended to improve consistency within the overall dataset, including positioning themselves centrally within their web camera frame, sitting in a well lit room, not allowing other faces or people to be present in the frame, and minimizing background noise.
For scripted responses (see Appendix A), participants were asked to record themselves repeating a short script while looking into the camera. Scripted responses were obtained using transcripts of the TIMIT dataset [8]. The TIMIT dataset consists of 462 real female and male American-English speakers, uttering a total of 1,718 short-to-medium length phonetically-rich sentences. Sentences of length less than the mean of 50 characters were removed. Ten sentences were then selected at random for the standardized scripts read by all participants. The remaining 728 sentences comprised the randomized scripts, for which each participant read a random sample of 10.
Participants then responded to four open-ended unscripted questions and asked to aim for a response that was close to 30 seconds in length. These were followed by quick-fire questions in which they repeated the question and provided a short response. Following the scripted and unscripted responses, participants were asked to perform simple actions using head and hand gestures. See Appendix A for the full list of prompts and scripts used.
We manually removed any invalid responses from the final dataset. This included ensuring that only a single person was visible in the video, the scene was reasonably well lit, and each submitted video clip contained a valid audio and visual track.
2.1 Version 1.0
We collected data from 220 participants for version 1.1 (version 1.0 had only 219 participants due to a corrupt file). The demographics of the participants are as follows (some participants identified with more than one race/ethnicity):
@@
  • Gender: 110 male, 106 female, 4 non-binary
  • Race/Ethnicity: 159 White/Caucasian, 38 Black/African American, 25 Asian, 3 American Indian/Alaska Native, 1 Native Hawaiian/Other Pacific Islander, 5 other.
Participants were asked to perform four simple visual actions: (1) wave their hand in front of their face while counting 1,2,3; (2) look down and right, straight down, and down and left, each time holding and counting 1,2,3; (3) look up and right, straight up, and up and left, each time holding and counting 1,2,3; (4) lean towards the camera without counting.
The audio/video recordings were converted from their initial .webm format to .mp4 and organized by prompt type using the H.264 codec with the libx264 library (this final video conversion is not performed in version 2.0 in order to minimize re-encoding).
Because version 1.1 corrects errors from version 1.0, we recommend using this version.
read wave hand turn head right/up tilt head down lean in
Figure 1: Representative examples of different prompt types (see Appendix A).
2.2 Version 2.0
We collected data from 280 participants for version 2.0. The demographics of the participants were as follows (some participants identified with more than one race/ethnicity):
  • Age: Range = 18-73 years, Mean = 38 years; standard deviation = 11.6 years
  • Gender: 146 male, 129 female, 3 non-binary, 2 not provided
@@
  • Race/Ethnicity: 203 White/Caucasian, 49 Black/African American, 20 Asian, 11 American Indian/Alaska Native, 1 Native Hawaiian/Other Pacific Islander, 10 other, 1 prefer not to say.
Participants were asked to perform the same four simple visual actions as in version 1.0: (1) wave their hand in front of their face while counting 1,2,3; (2) look down and right, straight down, and down and left, each time holding and counting 1,2,3; (3) look up and right, straight up, and up and left, each time holding and counting 1,2,3; (4) lean towards the camera without counting). An additional set of three new actions were added to this version: (5) pretending to yawn; (6) pretending to laugh for between 2 to 5 seconds; (7) clapping loudly three times, pausing for about 5 seconds between claps.
While audio/video recordings for version 1.0 were converted from their initial .webm format to .mp4 using the H.264 codec with the libx264 library, an FFmpeg “copy” command was used in version 2.0 without re-encoding to preserve recording quality.
@@ 3 Audio Generation
Participants were first asked to record themselves reading 10 consecutive phonetically-rich sentences, sourced from List 1 of the standard Harvard Sentences [27]. Participants were then asked to repeat the standard elicitation paragraph from the Speech Accent Archive, a phonetically comprehensive passage comprising a breadth of vowels and consonants [33]. These two scripted responses were used for the purpose of voice cloning, and had an average length of 30 seconds.
Using each participant’s cloned voice, a synthetic audio was created in their voice saying the same thing as in the original audio/video. For the unscripted responses, the original audio was transcribed using OpenAI’s Whisper, and for the scripted responses, we assumed that the participant correctly read the script. These text transcriptions were then provided to each voice cloning generators’ API to generate matching synthetic voices. The following generators were used for each version of the DeepSpeak dataset.
3.1 Version 1.0
Voice clones were generated using the Elevenlabs Create Voice endpoint [^3] https://elevenlabs.io/docs/api-reference/voices/add. The Create Speech API [^4] https://elevenlabs.io/docs/api-reference/text-to-speech/convert was used for speech synthesis. The eleven_multilingual_v2 model was used, with audio output returned in MP3 format at 44.1 kHz. The API was accessed in April 2024.
3.2 Version 2.0
As in version 1.0, voice clones in version 2.0 used the same two scripted prompts for all voice-cloning generators.
The ElevenLabs API was accessed as described in the previous section.
Voice clones were generated using the PlayAI Instant Voice Cloning endpoint [^5] https://docs.play.ht/reference/api-create-instant-voice-clone. The pyht Python package, with TTSOptions set to default parameters was used for speech synthesis, with audio output provided in MP3 format, at a sampling rate of 24 kHz.
Voice clones were generated using the Speechify v1 Voices API [^6] https://docs.sws.speechify.com/v1/api-reference/api-reference/tts/voices/create. The Speech API endpoint [^7] https://docs.sws.speechify.com/v1/api-reference/api-reference/tts/audio/speech was used for speech synthesis. The simba-english model was used, with output audio returned in WAV format at 48 kHz.
All API calls were made in November 2024.
@@ 4 Video Generation
We generated three types of video deepfakes: face swap, lip sync, and avatar, each of which is described next. Between version 1.0 and 2.0 of DeepSpeak, deepfake generators improved considerably. Because we are primarily interested in creating a dataset using state-of-the-art generators, we do not, in version 2.0, re-use the same generators as in version 1.0.
All source videos were re-encoded using FFmpeg with the H.264 codec at a constant rate factor of 18. The encoding present was set to slow, and the original audio stream was preserved without re-encoding.
4.1 Face-Swap
Face-swap deepfakes are created by replacing – eyebrow to chin and cheek to cheek – the original identity in a video with a new identity. In order to create a visually compelling face-swap deepfake, each identity in the full dataset is paired with a perceptually similar face as follows. Each identity is first represented by a mean CLIP embedding [22] of five randomly extracted video frames. For each identity a unique matched identity is assigned using the agglomerative clustering algorithm with cosine distance and cluster size constraint from the scikit-learn library [^8] https://scikit-learn.org/stable/modules/generated/sklearn.cluster.AgglomerativeClustering.html. Representative examples of these pairs are shown in Figures 2 and 3.
@@ 4.1 Face-Swap
After pairing each identity, the video of the original identity and a single frame of the matched identity are provided to the face-swap synthesis engine. The single frame is initially chosen to be the fifth frame in a randomly selected video of the matched identity. We found that if the eyes are closed in the matched face, the resulting face-swap deepfake suffered in quality. As such, we used MediaPipe [18] to extract facial features and ensured that the distance between the top and bottom eyelid landmarks was greater than a specified threshold. If this constraint failed, then the tenth frame was selected for consideration; this process was repeated, skipping five frames each time, until a suitable frame was found.
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 2: Four representative examples of matched identities from version 1.0.
4.1.1 Version 1.0
We used three variants of the open-source library FaceFusion as follows:
  • FaceFusion Live: The third face-swap configuration wraps the same FaceFusion configuration in a simulated live input streaming environment. Because our hardware (a Tesla T4 GPU) cannot generate video frames in real-time ($2424$ to $3030$ frames per second (fps)), the effective frame-rate of the generated video is decreased to approximately $1212$ fps.
4.1.2 Version 2.0
Between the construction of version 1.0 and version 2.0, FaceFusion expanded to include a larger suite of face-swap generators. For version 2.0, we used four variants from the FaceFusion library:
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 3: Four representative examples of matched identities from version 2.0.
4.2 Lip-Sync
Whereas the face-swap deepfake (Section 4.1) replaces an entire face with a new identity, a lip-sync deepfake modifies the mouth region to be consistent with a different audio track.
@@ 4.2 Lip-Sync
Given an original video and associated audio, we create two types of lip-sync deepfakes: (1) a lip-sync deepfake with an audio of the same identity extracted from a different video (i.e., the audio and video are now mismatched); and (2) a lip-sync deepfake with an AI-generated voice of the same identity (Section 3) with a transcript taken from a different video.
4.2.1 Version 1.0
Two types of lip-sync deepfakes were created using the following models:
4.2.2 Version 2.0
Two types of lip-sync deepfakes were created using the following models:
4.3 Avatar
Avatar deepfakes animate the head and lip movements of a static image to match a target video or audio track. Unlike face-swap and lip-sync deepfakes, which modify an existing video, avatar deepfakes generate movement from a single static image. Version 1.0 contains no avatar deepfakes because these models were still in their infancy.
4.3.1 Version 2.0
Three types of avatar deepfakes were created using the following models:

### v4

@@ Abstract
Deepfakes represent a growing concern across domains such as impostor hiring, fraud, and disinformation. Despite significant efforts to develop robust detection classifiers to distinguish the real from the fake, commonly used training datasets remain inadequate: relying on low-quality and outdated deepfake generators, consisting of content scraped from online repositories without participant consent, lacking in multimodal coverage, and rarely employing identity-matching protocols to ensure realistic fakes. To overcome these limitations, we present the DeepSpeak dataset, a diverse and multimodal dataset comprising over 100 hours of authentic and deepfake audiovisual content. We contribute: i) more than 50 hours of real, self-recorded data collected from 500 diverse and consenting participants using a custom-built data collection tool, ii) more than 50 hours of state-of-the-art audio and visual deepfakes generated using 14 video synthesis engines and three voice cloning engines, and iii) an embedding-based, identity-matching approach to ensure the creation of convincing, high-quality identity swaps that realistically simulate adversarial deepfake attacks. We also perform large-scale evaluations of state-of-the-art deepfake detectors and show that, without retraining, these detectors fail to generalize to the DeepSpeak dataset. These evaluations highlight the importance of a large and diverse dataset containing deepfakes from the latest generative-AI tools.
Data & Code: https://github.com/hfaridlab/deepspeak
Note that the DeepSpeak dataset is released in versions. See Appendix A for usage and release information.
@@ 1 Introduction
Today, generative-AI is capable of creating hyper-realistic images nightingale2022ai , voices barrington2025people , and videos groh2022deepfake of people talking or doing just about anything. These technologies hold the promise to both revolutionize many industries while also amplifying the spread and belief in dangerous lies and conspiracies chesney2019deep ; vaccari2020deepfakes , interfering with elections ferrara2024charting ; stockwell2024ai , super-charging small- and large-scale fraud bateman2022deepfakes , and – seemingly unable to escape its roots – continue to be used in the creation of non-consensual intimate imagery (NCII) and child sexual abuse material (CSAM) ding2025malicious ; viola2023designed .
Scalable, generalizable, and accurate detection of deepfakes has, therefore, become a pressing problem with deep social, political, and economic implications. At the same time, the nascent digital forensics community has struggled with the lack of large-scale, high-quality, up-to-date, and ethically collected datasets for training and evaluation.
In this work, we introduce an audio and video dataset designed to aid the digital-forensic, computer-vision, and broader AI-safety communities. This dataset consists of 100 hours of real and deepfake video of people talking and gesturing. The real videos were recorded with consent from the participants. The deepfakes consist of avatar (from three generators), face-swap deepfakes (with multiple variants from three generators), lip-sync deepfakes (from four generators), and audio deepfakes (from three generators) spliced into a subset of the lip-sync deepfakes. Table 1 presents a comparison between our dataset and recent datasets released over the past seven years.
We focus exclusively on “talking heads” in which one person is talking or gesturing in a typical video conferencing setup. We focus on this context because of the distinct harms that have emerged from both offline and real-time deepfake impersonations (including impostor hiring, fraud, and disinformation) and because existing datasets in this domain do not reflect the current quality and diversity of deepfake generators, while bearing ethical, practical, and legal shortcomings (see Table 1). Specifically, existing datasets largely comprise low-quality, outdated deepfake generators, where underlying data was scraped without participant consent. Moreover, these datasets do not include all types of deepfake generators and attack settings. A more comprehensive review of other datasets is included in Appendix B.
To remedy these shortcomings, our work makes the following contributions:
  • Documentation and Release of DeepSpeak. We introduce a methodology for the large-scale collection of real video recordings, self-submitted by a diverse selection of consenting participants (Section 2), along with the procedures used to generate corresponding deepfake video and audio (Sections 4 and 5). The dataset is publicly available by request under both research and commercial licenses.
  • Data Collection Tool. We provide a codebase for a web-based application designed to facilitate participant-led remote data collection. The tool supports the recording, storage and organization of webcam footage submitted by participants. When used in conjunction with our collection survey, the collected data is phonetically rich and diverse in terms of speech content, video durations, gestures, and includes both scripted and unscripted segments.
  • Method for Identity Matching. We devise a method for matching participants based on their visual and vocal features to create more convincing face-swap deepfakes, consistent with real-world deepfake attacks (Section 3). These methods are fully open-sourced.
  • Large-scale Benchmarking and Generalization Study. We perform large-scale evaluations of state-of-the-art deepfake detectors and show that these detectors, trained on others datasets (Table 1) fail to accurately distinguish between real and fake audio and video in DeepSpeak (Section 6). These evaluations highlight the importance of a large and diverse dataset containing deepfakes from the latest generative-AI tools.
Figure 1: An overview of the DeepSpeak Dataset sourced from a diverse selection of consenting participants using a custom-built data collection methodology. The dataset also comprises deepfakes generated from 14 video and three audio deepfake methods using facial identity matching to improve the realism of the generated deepfakes.
@@
FaceForensics  rossler2018faceforensics 2018 NA 1,004 N $\checkmark$ - - - 2,008
FaceForensics++ rossler2019faceforensics 2019 NA 1,000 N $\checkmark$ - $\checkmark$ - 4,000
DFDC dolhansky2020deepfake 2020 3,426 23,654 Y $\checkmark$ - - - 104,500
Celeb-DF li2020celeb 2020 59 590 N $\checkmark$ - - - 5,639
WildDeepfake zi2020wilddeepfake 2020 707 3,805 N NA NA NA NA 3,509
DeeperForensics jiang2020deeperforensics 2020 100 50,000 Y $\checkmark$ - - - 10,000
FakeAVCeleb khalid2021fakeavceleb 2021 600 570 N $\checkmark$ $\checkmark$ - - 25,000
ForgeryNet he2021forgerynet 2021 5,400 99,630 Y/N $\checkmark$ - - - 121,617
LAV-DF cai2022you 2022 153 36,431 N - - $\checkmark$ - 99,873
DF40 yan2024df40 2024 NA NA N - 100,000+
DeepSpeak (Ours) 2025 500 16,043 Y $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 14,005
@@ 2 Data Collection
The data collection was performed in four steps. Data collection for DeepSpeak was determined to qualify for exempt status by UC Berkeley’s Office for Protection of Human Subjects (OPHS).
Participant Recruitment.
Participants were crowd-sourced through the Prolific research recruitment platform. Participants were asked to give their consent for including their recordings, without any other identifying information, in a public dataset. Details of the consent statement can be found in Appendix P. A total of 500 participants were selected from a stratified sample ensuring equal distribution of gender, and with all participants reported as being native English speakers and U.S. residents, with demographics as follows (some participants identified with more than one race/ethnicity):
@@
  • Gender: 256 male, 235 female, 7 non-binary, 2 not provided
@@
  • Race/Ethnicity: 362 White/Caucasian, 87 Black/African American, 45 Asian, 14 American Indian/Alaska Native, 2 Native Hawaiian/Other Pacific Islander, 15 other, 1 prefer not to say.
Survey.
The data collection survey was designed to capture both speech and visual actions. For speech, it included phonetically rich audio data spanning varied audio durations with both scripted vs. conversational-style responses. Each participant was instructed to record themselves responding to between 32 and 35 separate prompts. Participants were paid \$7 for their time. The first two prompts were used for voice-clone training data (see Section 4). The remaining prompts were divided into four categories: (1) 10 standardized scripted responses in which each participant read the same prompt; (2) 10 randomized scripted responses in which participants read a randomized prompt; (3) 10 unscripted responses in which participants responded to questions; and (4) between 5-8 actions in which participants performed simple actions. Scripted responses were generated using transcripts of the TIMIT dataset, a linguistics research dataset consisting of utterances from 462 real female and male American-English speakers. See Appendix Q for the full list of prompts and scripts used.
Data Collection tool.
Both audio and video were recorded using a custom-built Google Chrome web application. The JavaScript and Python repository for this web application is available at https://github.com/hfaridlab/deepspeak/tree/main/data_collection. Details of the encoding and data pre-processing associated with the tool can be found in the Appendix C.
Validation.
Participants were given written and visual instructions to allow them to practice recording themselves and test their hardware. Participants were asked to adhere to a series of recording conditions intended to improve consistency within the overall dataset. We manually removed any invalid responses from the final dataset that did not meet these requirements. The details of this can be found in the Appendix C
3 Identity Matching
During manual inspection of the collected data, we observed that, albeit diverse in age, gender, and ethnicity, our collected data contains many individuals with similar facial and vocal features. In order to exploit this feature of the dataset and create more compelling deepfakes, each identity in the dataset was paired with another, perceptually and acoustically similar one. The code for producing this visual and audio matching, as well as the resulting visual and audio pairs is open-sourced at https://github.com/hfaridlab/deepspeak/tree/main/identity_matching.
Visual Matching.
Each identity is first represented by the average CLIP embedding [^1] https://github.com/OpenAI/CLIP radford2021learning extracted from five random video frames (filtered for low-quality frames, see Section 5.2). Shown in Figure 2 is a t-SNE visualization of a subset of these embeddings. Comparing this representation against the self-reported demographic information reveals that these CLIP embeddings cluster based on gender, ethnicity and facial similarity.
Figure 2: A t-SNE visualization of CLIP embeddings from real participant’s videos. The four highlighted pairs correspond to identities with maximal similarity as measured by the cosine distance between CLIP embeddings. Perceptually similar identities cluster in this t-SNE representation. The red/blue color coding corresponds to people who identify as female/male, which also clusters in this t-SNE representation.
For each identity, a unique matched identity is assigned using the agglomerative clustering algorithm with cosine distance and cluster size constraint from the scikit-learn library [^2] https://scikit-learn.org/stable/modules/generated/sklearn.cluster.AgglomerativeClustering.html. Additional examples of visual pairs are shown in Appendix I.
Audio Matching.
While the DeepSpeak dataset only matches a real voice to its own clone, for completeness, we describe a similar approach for audio identity matching. A single audio from a given vocal identity is represented by a 192-dimensional embedding generated by TitaNet-L, a neural model primarily used for speaker recognition. The same short scripted audio was used for all identities. Each identity can then be matched using the same process outlined for visual identity matching. Further details of this approach are provided in Appendix J.1.
@@ 4 Audio Generation
Participants were first asked to record themselves reading 10 consecutive phonetically-rich sentences, sourced from List 1 of the standard Harvard Sentences rothauser1969ieee , a collection of sentences representing best practice for standardized evaluation of speech processing and audio quality in controlled settings. Participants were then asked to repeat the standard elicitation paragraph from the Speech Accent Archive, a phonetically comprehensive passage comprising a breadth of vowels and consonants saa . These two scripted responses were used for the purpose of voice cloning, and had an average length of 30 seconds.
Using each participant’s cloned voice, a synthetic audio was created in their voice saying the same thing as in the original audio/video. For the unscripted responses, the original audio was transcribed using OpenAI’s Whisper, and for the scripted responses, we assumed that the participant correctly read the script. These text transcriptions were then provided to each voice cloning generators’ API to generate matching synthetic voices.
Voice clones were generated using three commercial cloning and Text-to-Speech (TTS) services: ElevenLabs, PlayAI and Speechify. The details of API end points used, alongside parameters, can be found in Appendix E.
@@ 5 Video Generation
We generated three types of video deepfakes: face swap, lip sync, and avatar, each of which is described next. The resulting dataset is randomly split into 80/20 training/testing splits with no overlap in facial or voice identities. A breakdown of the resulting dataset’s statistics, including the total file size (GB), file counts (N), and video length (hrs) are included in Appendix D.
Figure 3: Overview of DeepSpeak deepfakes: Face-swap deepfakes replace the facial region from a source identity with the face of a target identity while retaining the audio from the target video. Lip-sync deepfakes overlay a generated mouth region onto the original face and synchronize it with a new audio track (real or fake). Avatar deepfakes animate the head and shoulders of an identity based on a single still frame and a new audio (real or fake).
5.1 Generation
Face-Swap.
Face-swap deepfakes are created by replacing – eyebrow to chin and cheek to cheek – the original identity in a video with a new identity. We swapped faces of identity pairs identified through the visual matching (see Section 3). This ensured that the swapped identities were perceptually similar to begin with, which made for more compelling deepfakes. This resembles conventional practices of in-the-wild deepfake production, where actors are chosen based on their similarity to the target identity.
@@ 5.1 Generation
An overview of face-swap deepfake generation is shown in Figure 3, row one. To generate a face swap, the video of the original identity and a single frame of the matched identity are provided to the face-swap synthesis engine. The single frame is initially chosen to be the fifth frame in a randomly selected video of the matched identity. We found that if the eyes are closed in the matched face, the resulting face-swap deepfake suffered in quality. As such, we used MediaPipe lugaresi2019mediapipe to extract facial features and ensured that the distance between the top and bottom eyelid landmarks was greater than a specified threshold. If this constraint failed, then the tenth frame was selected for consideration; this process was repeated, skipping five frames each time, until a suitable frame was found. We used seven face-swap methods, as detailed in Appendix D.
Lip-Sync.
@@
Whereas the face-swap deepfake replaces an entire face with a new identity, a lip-sync deepfake modifies the mouth region to be consistent with a different audio track. An overview of lip-sync deepfake generation is shown in row two of Figure 3. Given an original video and associated audio, we create two types of lip-sync deepfakes: (1) a lip-sync deepfake with an audio of the same identity extracted from a different video (i.e., the audio and video are now mismatched); and (2) a lip-sync deepfake with an AI-generated voice of the same identity (Section 4) with a transcript taken from a different video. Four methods of lip-sync deepfakes were employed, as further described in Appendix D.
Avatar.